# Multi-state Protein Design with DynamicMPNN

### v1

@@ 4 Conclusion
We present DynamicMPNN, the first explicit multi-state inverse folding model, achieving up to 13% improvement over ProteinMPNN on our multi-state benchmark. By jointly learning across conformational ensembles rather than aggregating single-state predictions, DynamicMPNN captures sequence constraints required for multiple functional conformations. This opens possibilities for engineering synthetic bioswitches, allosteric regulators, and molecular machines.
Software and Data
We provide the code for our model in the following repository: github.com/Alex-Abrudan/DynamicMPNN.
Acknowledgements
The authors would like to thank Adam Broerman, Rohith Krishna, Kai Yi, Simon Mathis, Hannah Han, Vladimir Radenkovic, Gustavo Parisi, Sebastian Ahnert, and Pietro Liò for insightful comments and discussions. CKJ was supported by the A*STAR Singapore National Science Scholarship (PhD). AK was supported by the Graduate School of Quantitative Biosciences Munich (QBM).
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
Appendix A Supplementary Results
A.1 Additional Performance Analysis
For our statistical analysis, we first evaluated the normality of the paired differences between model performances (RMSD values) using the Shapiro-Wilk test. The test results (W = 0.913, p 0.001) strongly rejected the null hypothesis of normality for the distribution of differences.
Therefore, we selected the Wilcoxon signed-rank test as the appropriate non-parametric alternative to the paired t-test for assessing statistical significance between model performances. This test makes no assumptions about the underlying distribution shape, making it robust for our analysis of paired protein structure comparisons across the different design approaches.
Fig. LABEL:fig:supplementary_plots (a), (e) shows how DynamicMPNN maintains its superiority for the single-state best sequence design for the RMSD and pLDDT metrics. However, averaging the performance of all designed sequences erases the advantage of DynamicMPNN over ProteinMPNN and natural sequences (Fig. LABEL:fig:supplementary_plots (b), (c), (d), (f)).
While ProteinMPNN outperforms DynamicMPNN on sequence recovery and perplexity (Table 2), Wang et al. (2023) argue that refoldability - emphasised in the present work - is more direct, structure-grounded, and superior to sequence recovery and perplexity when it comes to evaluating inverse folding models.
| Model | Sequence Recovery | Perplexity |
| DynamicMPNN | 0.29 | 10.2 |
| ProteinMPNN MSD | 0.38 | 5.12 |
Appendix B Dataset Details
@@
To construct the dataset, we obtained 46,033 Multiple Sequence Alignment (MSA) clusters at local sequence similarity from the latest version of CoDNaS (v2025) (Monzon et al., 2016), including NMR model structures. Importantly, the CoDNaS dataset prevents potential errors created by mixing different homologues in the same cluster by enforcing the same UniProt ID for all custer members. The 95% similarity threshold accounts for soluble tags or point mutations/alterations performed in different experiments so that all available conformations of a protein are included - as different experimental conditions and sequence variations can reveal distinct thermodynamic states of the same protein (Best et al., 2006). While clusters contain varying numbers of conformations (Fig. 2b) we constructed our dataset using only pairs of chains from one or two PDB entries that have the largest RMSD from each cluster (Fig. 2c). Since the proteins targeted by this project explore two functional states, this choice aims to minimize alignment artifacts and maximize conformational signal during training - chosen pairs represent the most distinct conformational states. We will include conformational information of the whole cluster in future work.
Appendix C Multi-conformational inverse folding
Given a set of target conformations we seek to model the conditional probability distribution over amino acid sequences that can adopt all specified structures. The challenge lies in learning a sequence distribution that simultaneously satisfies multiple structural constraints. We decompose this joint conditional probability using the autoregressive factorization:
@@ D.4 Conformation order-invariant pooling
After using message passing layers that are conformation order-equivariant, we add a conformation order-invariant head, which performs average pooling across the conformation channel of the scalar and vector feature tensors, similar to Joshi et al. (2025): and to and , where is the sequence length, is the number of backbones, is the number of scalar features, and is the number of vector features. The only pooling strategy used in this work is the pooling of the maximum RMSD pair of chains - therefore - although more pooling strategies for homo-oligomers can be used, such as equal averaging of all chains to be inverse folded in the selected PDB entries.
Appendix E Further Discussion
E.1 Motivation
A commonly derived assumption from Anfinsen’s experiment is that proteins adopt only one native 3D structure, leading to the “one sequence, one structure, one function” canon. This view has been indirectly reinforced by the predominant use of X-Ray crystallography in experimental protein structure determination - it requires the protein to form a single, diffractible crystal (Dishman & Volkman, 2018). This led to a bias in the Protein Data Bank towards globular, thermodynamically stable chains. However, the dynamic nature of protein NMR and the advancements of Cryo-EM as well as the redundancy among PDB entries (i.e. multiple entries that contain almost identical protein sequences) have offered valuable insights into protein conformational landscape (Hrabe et al., 2015; Monzon et al., 2016). A series of MD datasets simulating diverse protein structures have been released (Vander Meersche et al., 2023; Mirarchi et al., 2024), albeit they have limited sequence variability and only a fraction of the simulations provide valuable information into dynamics of large molecular conformational changes.
The large number of high-resolution deposited structures in the PDB allowed for the training of large geometric deep learning models that boosted the prediction accuracy of protein structure (Jumper et al., 2021; Kryshtafovych et al., 2019). Remarkable breakthroughs in protein backbone inverse folding (Dauparas et al., 2022) and de novo backbone generation have followed suit (Watson et al., 2023). ProteinMPNN (Dauparas et al., 2022) in particular has streamlined single-state protein inverse folding due to much smaller computational and time costs coupled with increased sequence recovery, when compared to physics-based methods (Liu & Kuhlman, 2006) that treat sequence design as an energy optimization problem. While these models have been successfully used for de novo protein design for improved function (Sumida et al., 2024), they mostly rely on enhancing protein stability and solubility rather than dealing with dynamics and interaction mechanisms.

### v2

@@ 4 Conclusion
We present DynamicMPNN, the first explicit multi-state inverse folding model, achieving up to 25% improvement over ProteinMPNN on our multi-state benchmark. By jointly learning across conformational ensembles rather than aggregating single-state predictions, DynamicMPNN captures sequence constraints required for multiple functional conformations. This opens possibilities for engineering synthetic bioswitches, allosteric regulators, and molecular machines. It is unclear if the presented one-size-fits-all approach to multi-state design will be effective experimentally, or if specialized models for different classes of conformational changes depending on their thermodynamic complexity will be beneficial.
Appendix A Appendix
A.1 LLM acknowledgement
The authors acknowledge that they have used LLMs in the process of paper writing for style suggestions and proofreading.
Appendix B Supplementary Results
B.1 DynamicMPNN preliminary designs
Appendix C Dataset Details
@@
To construct the dataset, we obtained 46,033 Multiple Sequence Alignment (MSA) clusters at local sequence similarity from the latest version of CoDNaS (v2025) (Monzon et al., 2016), including NMR model structures. Importantly, the CoDNaS dataset prevents potential errors created by mixing different homologues in the same cluster by enforcing the same UniProt ID for all custer members. All available conformations of a protein are included - as different experimental conditions and sequence variations can reveal distinct thermodynamic states of the same protein (Best et al., 2006). While clusters contain varying numbers of conformations (Fig. 2b) we constructed our dataset using only pairs of chains from one or two PDB entries that have the largest RMSD from each cluster (Fig. 2c). Chosen pairs represent the most distinct conformational states. We will include conformational information of the whole cluster in future work.
While other inverse folding models (Hsu et al., 2022) saw improved performance when trained on orders of magnitude of more protein structures from AFDB, previous studies have found that the majority of high-confidence structures in AFDB map to known CATH superfamilies (Bordin et al., 2023), and that AlphaFold struggles in predicting alternative states (Chakravarty et al., 2024). We therefore decided against including AFDB structures in our training set.
@@ D.4 Conformation order-invariant pooling
After using message passing layers that are conformation order-equivariant, we add a conformation order-invariant head, which performs average pooling across the conformation channel of the scalar and vector feature tensors, similar to Joshi et al. (2025) (2025): and to and , where is the sequence length, is the number of backbones, is the number of scalar features, and is the number of vector features. The only pooling strategy used in this work is the pooling of the maximum RMSD pair of chains - therefore - although more pooling strategies for homo-oligomers can be used, such as equal averaging of all chains to be inverse folded in the selected PDB entries.